OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

OpenStudy (anonymous):

m = y2 -y1 / x2 - x1. y = mx + b (-3,2) = (x1, y1) (2,1) = (x2,y2) \[m = \frac{ 1-2 }{ 2 - -3 }\] m = -1/ 5 y = (-1/5)x +b plug in either (-3,2) or 2,1) to solve b